A 5 1/3 ounce can of olives sell for $1.69. At this rate, how much should an 8 ounce can coast . I need to write a proportion and then find the answer
Answer by macston(5194)

$1.69/5.33oz=x/8.00oz Multiply each side by (5.33oz)(8.00oz)
$1.69(8.00oz)=x(5.33oz) Divide each side by (5.33oz)
($1.69*8.00oz)/5.33oz=x
$2.54=x ANSWER: An 8 ounce can would cost $2.54.